Mark and Eduardo are selling pies for a school fundraiser. Customers can buy cherry pie and lemon meringue pies. Mark sold 8 cherry pies and 6 lemon meringue pies for a total of 192. Eduardo sold 8 cherry pies and 7 lemon meringue pies for a total of 212. Find the cost of each one cherry pie and one lemon meringue
Answer by VFBundy(438)

Cost of one cherry pie = C
Cost of one lemon meringue pie = L

8C + 6L = 192
8C + 7L = 212

Multiply first equation by -1:

-8C - 6L = -192
8C + 7L = 212

Add both equations together:

L = 20

We now know L = 20. Plug this into either original equation. (I chose the first one.):

8C + 6L = 192

8C + 6(20) = 192

8C + 120 = 192

8C = 72

C = 9

Cost of one cherry pie = C = 9
Cost of one lemon meringue pie = L = 20